Bellingham Company produces a product that requires 12 standard pounds per unit. The standard price is $9.5 per pound. If 3,400 units used 42,400 pounds, which were purchased at $9.02 per pound, what is the direct materials (a) price variance, (b) quantity variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

  1. Direct Materials Variances

    Bellingham Company produces a product that requires 12 standard pounds per unit. The standard price is $9.5 per pound. If 3,400 units used 42,400 pounds, which were purchased at $9.02 per pound, what is the direct materials (a) price variance, (b) quantity variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

    a. Direct materials price variance $ Favorable
    b. Direct materials quantity variance $ Unfavorable
    c. Direct materials cost variance $ Favorable
Answer #1

Direct material price variance = (Standard price-actual price)actual qty = (9.5-9.02)*42400 = 20352 Favorable

Direct material quantity variance = (Standard quantity-actual qty)Standard price = (3400*12-42400)*9.5 = 15200 Unfavorable

Direct material cost variance = (3400*12*9.5)-(42400*9.02) = 5152 Favorable